OpenStudy (vianne):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (anonymous):

Didn't get the question properly.. :(

OpenStudy (vianne):

:/ I'm not really sure what the logic is to solve this, I have 4 more questions.. ;-;

OpenStudy (vianne):

Why does it put ��������� in the brackets? O_o

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Okay..

OpenStudy (vianne):

Nevermind.. I'm guessing it's C and D because, (-1) + 1 = 0, but 0 isn't in the brackets.. And same goes for subtraction since it'll end up with -2. While Multiplication & Division have -1/1 as the product/quotient..

OpenStudy (anonymous):

Multiplication is okay.. :)

OpenStudy (anonymous):

Division is also looking good to me.. :)

OpenStudy (vianne):

I'll see if I get it correct then reply when I'm done with the other 4 questions. =D

OpenStudy (anonymous):

What are other 4?

OpenStudy (vianne):

Which equations show that the set of whole numbers is closed under multiplication? Choose all answers that are correct. A. –1 • –1 = 1 B. 0 • –1 = 0 C. 0 • 2 = 0 D. 2 • 1 = 2 This is the second question.

OpenStudy (anonymous):

Whole numbers are : 0, 1, 2, ..................

OpenStudy (vianne):

Negative numbers don't count, correct?

OpenStudy (anonymous):

Negative numbers are not allowed.. So, you are left with two options..

OpenStudy (vianne):

C And D?

OpenStudy (anonymous):

Good.. :)

OpenStudy (vianne):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (vianne):

That's the third question..

OpenStudy (anonymous):

D is for sure answer :P

OpenStudy (vianne):

A?

OpenStudy (anonymous):

8 + 5 = In the set or not?

OpenStudy (vianne):

Oh right.... :/

OpenStudy (anonymous):

Even integer set, when you add two even you are going to get another even.. :)

OpenStudy (anonymous):

\(2 + 4 = 6\), \(6\) is also even.. :)

OpenStudy (vianne):

Ohk..

OpenStudy (vianne):

Are rational numbers like a/b and can't end up with 0?

OpenStudy (anonymous):

Even Integer Set is : \(\{..... -4, -2, 0, 2, 4, 6...........\}\)

OpenStudy (anonymous):

I did not get you.. :(

OpenStudy (vianne):

Hold on, lemme google rational numbers. :3 Kinda forgot what it was.

OpenStudy (zarkon):

a/b where a and be are integers and b is not zero

OpenStudy (anonymous):

Rational numbers contain whole number + integers.. Decimals also allowed, terminating and recurring too..

OpenStudy (vianne):

So is 4.5 rational? because it is 9/2 I think..?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

See, 4.5, the decimal ends at 5, terminating one, it is rational..

OpenStudy (vianne):

Ok

OpenStudy (vianne):

So then C. Rational Numbers wouldn't be the answer?

OpenStudy (zarkon):

the sum of two rational numbers is a rational number

OpenStudy (anonymous):

He he he, I was about to ask Zarkon for confirmation.. :)

OpenStudy (zarkon):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

If addition of two numbers is giving you a rational number which will be in the set, then it is your answer.. :)

OpenStudy (vianne):

Ok

OpenStudy (anonymous):

three answers in this.. Now what is number 4 ??

OpenStudy (vianne):

Numbers p and q are whole numbers. Which statements are always true? Choose all answers that are correct. A. p + q is a whole number B. p – q is a whole number C. p • q is a whole number D. is a whole number

OpenStudy (vianne):

I'm thinking that it's A, B, C

OpenStudy (vianne):

Since we don't know what P and Q is, it might be 6 divided by 4 or something..

OpenStudy (vianne):

Plus I don't think that a sum would end up with a decimal when there is 2 whole numbers.. And same goes with subtraction.

OpenStudy (anonymous):

they are whole numbers.. :P

OpenStudy (anonymous):

\(2 - 5 = \) whole number or not?

OpenStudy (vianne):

Yes, but a whole number can still contain a decimal for the quotient..

OpenStudy (vianne):

Ahh wasn't thinking about that.. :P

OpenStudy (anonymous):

Addition is perfect..

OpenStudy (vianne):

And so is multiplication?

OpenStudy (anonymous):

yep, good..

OpenStudy (vianne):

Should we head onto question 5?

OpenStudy (anonymous):

Division and subtraction can be killed here.. :)

OpenStudy (vianne):

Ohk.

OpenStudy (anonymous):

Any doubt anywhere?

OpenStudy (vianne):

Nope.

OpenStudy (anonymous):

For first one, division and multiplication hold good.. :)

OpenStudy (vianne):

Yep.

OpenStudy (anonymous):

Okay, good, you interacted well, I appreciate you for that.. :)

OpenStudy (vianne):

Question 5? :P

OpenStudy (anonymous):

Still?

OpenStudy (vianne):

We only did up to question 4. :P

OpenStudy (anonymous):

Or you have 4 without counting first one?

OpenStudy (anonymous):

Okay, let it come.. :P

OpenStudy (vianne):

Is the set of multiples of 6 closed under subtraction? Explain why or provide a counterexample if not. A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12.

OpenStudy (anonymous):

What about b?

OpenStudy (anonymous):

\(0\) is also a multiple of \(6\)..

OpenStudy (vianne):

'cause of 2 x 3 = 6..

OpenStudy (anonymous):

wait..

OpenStudy (vianne):

I thought 6 x 0 = 0?

OpenStudy (anonymous):

What do you think? What this set is?? Can you write this set?

OpenStudy (anonymous):

Write this set? write the set for multiples of \(6\).. :)

OpenStudy (vianne):

Okay.\[6 \times 1\]?

OpenStudy (vianne):

It has to be all multiplication right?

OpenStudy (anonymous):

You are given a set, do you know what that set is?

OpenStudy (vianne):

the set they gave me was multiples of 6.. Or 2 6's..

OpenStudy (anonymous):

The set so called "Multiples of \(6\)" is the set: \[\{..........-18, -12, -6, 0, 6, 12, 18..........\}\]

OpenStudy (anonymous):

Multiples of \(6\) means number which are divisible by \(6\), giving \(0\) as remainder.. :)

OpenStudy (anonymous):

getting?

OpenStudy (vianne):

A little bit

OpenStudy (anonymous):

2, 3 are factors of 6 but not the multiples of 6.. :)

OpenStudy (anonymous):

do you know the table for 6??

OpenStudy (vianne):

I do

OpenStudy (anonymous):

the numbers in the table are nothing but multiples of 6..

OpenStudy (vianne):

Yes.

OpenStudy (anonymous):